The square of one number is 23 less than the square of a second number. If the second number is 1 more than the first, what are the two numbers?
Answer by mananth(16949)

let one number be x
the other number be y
x^2=y^2-23
x^2-y^2=-23..............1
x+1=y
substitute the value of y in equation1
x^2-(x+1)^2=-23
x^2-x^-2x-1=-23
-2x=-23+1
-2x=-22
x=-22/-2
x=11
y= x+1
y=12